OpenStudy (anonymous):

Amy has 35 tiles of the same shape and size in a box. The colors of and amounts of the tiles are shown below: 12 red tiles 8 yellow tiles 15 pink tiles Without looking in the box, Amy takes out a tile at random. She then replaces the tile and takes out another tile from the box. What is the probability that Amy takes out a yellow tile in both draws? 8 over 35 multiplied by 7 over 34 equal 56 over 1190 8 over 35 multiplied by 8 over 35 equal 64 over 1225 8 over 35 plus 7 over 34 equal 517 over 1190 8 over 35 plus 8 over 35 equal 560 over 1225

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

we have 8 favorable outcomes and 12+8+15= 35 possible outcomes so probability= favorable outcomes/ possible outcomes=...

OpenStudy (michele_laino):

that is the probability associated to a single extraction

OpenStudy (michele_laino):

favorable outcomes = 8 possible outcomes = 35 \[\Large \frac{{{\text{favorable outcomes}}}}{{{\text{possible outcomes}}}} = ...\]

OpenStudy (anonymous):

so i have to divide

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

the probability is 8/35 it is simple!

OpenStudy (michele_laino):

for a single extraction

OpenStudy (anonymous):

so what about the other part

OpenStudy (michele_laino):

for second extraction, the probability is the same, namely it is 8/35, since Amy replaces the tile into the box. Now the requested probability is the product of the two probabilities, namely: probability = (8/35)*(8/35)=...?

OpenStudy (anonymous):

64/1225

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Yay!

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

A coin was flipped 150 times. The results of the experiment are shown in the following table: Heads Tails 90 60 Which of the following best describes the experimental probability of getting heads? It is 10% higher than the theoretical probability. It is 10% lower than the theoretical probability. It is equal to the theoretical probability for this data. The experimental probability cannot be concluded from the data in the table.

OpenStudy (michele_laino):

the theoretical probability is 1/2 since we have two possible outcomes, namely a head and a tail and we have one favorable outcome, namely a head, so probability = 1/2= 0.5 The experimental probability to get a head is 90/150=...

OpenStudy (michele_laino):

since we have 90 favorable cases and 150 possible cases or outcomes

OpenStudy (anonymous):

0.6

OpenStudy (anonymous):

SO IT 10% lower than the theoretical probability

OpenStudy (michele_laino):

now the difference between the experimental probability and the theoretical probability is: 0.6-0.5=...

OpenStudy (anonymous):

so its 10% higher

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

so the answer is A

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

yay!!!! last question sorry

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

The number of marbles of different colors stored in a hat is listed below: 8 red marbles 10 green marbles 6 blue marbles Without looking in the hat, Tessa takes out a marble at random. She replaces the marble and then takes out another marble from the hat. What is the probability that Tessa takes out a blue marble in both draws? fraction 1 over 16 fraction 1 over 12 fraction 1 over 4 fraction 1 over 2

OpenStudy (michele_laino):

here we have 8+10+6= 24 possible outcomes and we have 6 favorable outcomes so the probability associated to a single extraction is: p= 6/24 please simplify that fraction

OpenStudy (anonymous):

0.25

OpenStudy (michele_laino):

we can divide both numerator and denominator by 6, wha do you get?

OpenStudy (anonymous):

divide what?

OpenStudy (anonymous):

oh 1/4

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

so is that the answer @Michele_Laino

OpenStudy (michele_laino):

now, the probability associated to the second extraction is the same, namely 1/4. So the requested probability is given by the product between the above probabilities, namely p= (1/4)*(1/4)=...?

OpenStudy (michele_laino):

what is: \[\Large P = \frac{1}{4} \times \frac{1}{4} = ...?\]

OpenStudy (anonymous):

Dang it

OpenStudy (anonymous):

i just submitted my test

OpenStudy (anonymous):

and i got it wrong

OpenStudy (anonymous):

because i forgot to do the last part

OpenStudy (anonymous):

):

OpenStudy (anonymous):

but i got a 87% percent on my test

OpenStudy (michele_laino):

so have you passed it?

OpenStudy (anonymous):

yes i did

OpenStudy (michele_laino):

CONGRATULATIONS!!!!!

OpenStudy (anonymous):

im finished with math until next year yay!!!!!!!!

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

Im going to be in 8th grade

OpenStudy (michele_laino):

ok!! :)

OpenStudy (anonymous):

but Atleast i got a high B on my test !

OpenStudy (anonymous):

and i have a A in Math!

OpenStudy (michele_laino):

good job!!

OpenStudy (anonymous):

Thank you so much you really helped me undertand what i was doing. But i will be coming back just in case i need help lol!

OpenStudy (michele_laino):

ok! :)

OpenStudy (jack1):

hi marc just if u could, please only post 1x question per post cheers